please I need your help with this question, thanks in advance Capacitor Circuits C1=8uF А о — Ce=7uF - Cs=1MP - C₂=4pF во C3=8uIF What is the capacitance between terminals A and B (in mic
Answer #1

C = 8MF Hea Cq=74F -- Car - cq = 44f Bor - 3=8uf Here, (C2, C3) in series C > ( 2) ( 3) (8(4) 8+4 wloo B C + Ci - C23 - 4 - tc1235 + cq B C4, C235 - in parallel - CAB = C4 + 41235 = 7+2.514 [ SAB= 9.514 ME